# A geometric approach toconjugation-invariant random permutations

### v3

@@ 4.1 Decreasing subsequences
Suppose, without loss of generality, that $\delta=1/\beta$ for some integer $\beta$. Fix an arbitrary integer $K>16\beta^{2}$ and set $\Delta x:=1/K$. Slice $[0,1]^{2}$ into a regular grid $\left(C_{i,j}\right)_{1\leq i,j\leq K}$ made of square cells
@@
Notice that any down-right path of points can only visit at most one diagonal square cell $C_{i,i}$. Therefore
@@
where $\widetilde{\mathcal{P}}$ is the restriction of $\mathcal{P}$ to the cells $C_{i,j}$ with $i\neq j$. For each $i$, 3.1 asserts that $\mathcal{P}\cap C_{i,i}$ is a superposition of three (correlated) random-sized sets of i.i.d. uniform points in $C_{i,i}$. Writing $L_{m}$ for the $\mathrm{LDS}$ of a uniform permutation of size $m$, we obtain:
@@
for some constant $c_{\delta}>0$ which depends only on $\delta$ and whose value may change throughout the proof. Then, since $32\sqrt{2}K^{-1}<\sqrt{2}\delta^{2}/8<2\delta/$ 3, [DZ99, Theorem 2] implies
@@
for some other $c_{\delta}>0$. From Equations 3, 4 and 5 we deduce
@@
To study the remaining term, we can broadly rewrite the proof of [DZ95, Lemma 9]. Set $\Delta y:=\Delta x/\beta=1/(K\beta)$ and slice $[0,1]^{2}$ into a thinner grid $\left(\widetilde{C}_{i,j}\right)_{1\leq i\leq K,1\leq j\leq K\beta}$ made of rectangular cells
@@
Say a sequence of indices $(\mathbf{a},\mathbf{b})=(a_{i},b_{i})_{1\leq i\leq K}$ is admissible if it satisfies $1\leq b_{i+1}\leq a_{i+1}\leq b_{i}\leq a_{i}\leq K\beta$ for all $i\in[K-1]$. Informally, it encodes a down-right sequence of cells where $a_{i}$, resp. $b_{i}$, is the highest, resp. lowest, cell in the $i$-th column. See Figure 3 for a representation. Fix such $(\mathbf{a},\mathbf{b})$ and write $\widetilde{\mathcal{P}}_{(\mathbf{a},\mathbf{b})}$ for the restriction of $\widetilde{\mathcal{P}}$ to the cells associated with $(\mathbf{a},\mathbf{b})$. For each $1\leq i\leq K$, write
@@
Using Equations 7 and 8, we deduce:
@@
We can decompose the family $\widetilde{\mathcal{P}}\cap R_{i_{0}}=(\mathcal{P}\cap A)\sqcup(\mathcal{P}\cap B)$ into its points above and below $C_{i_{0},i_{0}}$. Since we have $\mathrm{Leb}(A)\leq\Delta x<\delta^{2}/16$, 3.2 yields:
@@
and likewise for $B$. 3.3 asserts that the sets $\mathcal{P}\cap A$ and $\mathcal{P}\cap B$ consist of i.i.d. uniform points, hence:
@@
The first term of (9) is handled in a similar way. If $i\neq i_{0}$, by 3.2:
@@
We can then use 3.3 and [DZ99, Theorem 2] as before:
@@
Finally, putting together Equations 9, 10 and 11:
@@
for some $c_{\delta}>0$, therefore by 3.3 and [DZ99, Theorem 1]:
@@
where $\delta>0$ is arbitrary. This concludes the proof of Theorem 2.1. ∎
Proof of 2.2.
We work conditionally given $t^{(n)}$. Then for each $n$, $\tau_{n}$ is a uniform $t^{(n)}$-cyclic permutation. Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$ and $\check{\mathcal{P}}$ be its points outside the diagonal. Define $\check{t}^{(n)}$ as the cycle type $(0,t_{2}^{(n)},t_{3}^{(n)},\dots)$ of size $\check{n}:=n-t_{1}^{(n)}$. Then according to 1.1, $\check{\tau}_{n}:={\rm Perm}\left(\check{\mathcal{P}}\right)$ is a uniform $\check{t}^{(n)}$-cyclic permutation. Since fixed points contribute at most $1$ to $\mathrm{LDS}(\tau_{n})$, we can write:
@@
Thus by Theorem 2.1, for any $\delta>0$ there exists a universal constant $c_{\delta}>0$ such that for any $n$:
@@
and likewise for the lower tail bound. If $\check{n}\to\infty$ in probability then the right-hand side goes to 00 as $n\to\infty$, and this proves convergence in probability. Convergence in $L^{q}$ follows as in the proof of Theorem 2.1, bounding $\mathrm{LDS}(\tau_{n})$ above by $n-t_{1}^{(n)}+1$ and using the fact that the previous tail inequalities hold a.s. conditionally given $t^{(n)}$. ∎
@@ 4.2 Increasing subsequences
Proof of 2.3.
Fix $\delta>0$. Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$ as in 1.1.
@@
Using 3.2, there exists $c_{\delta}>0$ such that:
@@
Then by [DZ99, Theorem 1] along with 3.3, denoting by $L_{m}$ the $\mathrm{LIS}$ of a uniform permutation of size $m$:
@@
since $\frac{1-\delta/2}{K}>\frac{1-\delta}{K-1}$. Finally, using Equation 12:
@@
For this, simply use 3.1 and a rough union bound. Since
@@
Proof of 2.4.
The method is the same as in the proof of 2.2, using rather the bounds
@@ 5 Proofs of the results on Robinson–Schensted shapes
Our proof of Theorem 2.7 relies heavily on the substantial work of Sjöstrand [Sjö23]. Let us recall a few definitions from the latter.
[…]
@@
Informally, $D(u)$ is the set of “south-west corners” in the integer level lines of $u$. If $u\in\mathcal{U}_{0,r}(\Omega)$ then $D(u)$ is an $\lfloor r\rfloor$-decreasing subset of $\Omega$ [Sjö23, Lemma 6.2]. Reciprocally, for any finite $k$-decreasing subset $P$ of $\Omega$, there exists a $\kappa_{P}\in\mathcal{U}_{0,k}(\Omega)$ such that $D(\kappa_{P})=P$ [Sjö23, Definition 3.7 and Lemma 6.2]. Therefore, if $\mathcal{P}$ is a finite subset of $\Omega$:

### v4

@@ 4.1 Decreasing subsequences
Suppose, without loss of generality, that $\delta=1/\beta$ for some integer $\beta$. Fix an arbitrary integer $K>16\beta^{2}$ and set $\Delta x:=1/K$. Slice $0,1]^{2}$ into a regular grid $\left(C_{i,j}\right)_{1\leq i,j\leq K}$ made of square cells
@@
Notice that any down-right path of points can only visit at most one diagonal square cell $C_{i,i}$. Therefore:
@@
where $\widetilde{\mathcal{P}}$ is the restriction of $\mathcal{P}$ to the cells $C_{i,j}$ with $i\neq j$. For each $i$, ˜3.1 asserts that $\mathcal{P}\cap C_{i,i}$ is a superposition of three (correlated) random-sized sets of i.i.d. uniform points in $C_{i,i}$. Writing $L_{m}$ for the $\mathrm{LDS}$ of a uniform permutation of size $m$, we obtain:
@@
for some constant $c_{\delta}>0$ which depends only on $\delta$ and whose value may change throughout the proof. Then, since $2\sqrt{2}K^{-1}<\sqrt{2}\delta^{2}/8<2\delta/3$, [S98, Theorem 2] implies
@@
for some other $c_{\delta}>0$. From Equations˜3, 4 and 5 we deduce
@@
To study the remaining term, we can broadly rewrite the proof of [DZ95, Lemma 9]. Set $\Delta y:=\Delta x/\beta=1/(K\beta)$ and slice $0,1]^{2}$ into a thinner grid $\left(\widetilde{C}_{i,j}\right)_{1\leq i\leq K,1\leq j\leq K\beta}$ made of rectangular cells
@@
Say a sequence of indices $(\mathbf{a},\mathbf{b})=(a_{i},b_{i})_{1\leq i\leq K}$ is admissible if it satisfies $1\leq b_{i+1}\leq a_{i+1}\leq b_{i}\leq a_{i}\leq K\beta$ for all $i\in[K-1]$. Informally, it encodes a down-right sequence of cells where $a_{i}$, resp. $b_{i}$, is the highest, resp. lowest, cell in the $i$-th column. See Figure˜3 for a representation. Fix such $(\mathbf{a},\mathbf{b})$ and write $\widetilde{\mathcal{P}}_{(\mathbf{a},\mathbf{b})}$ for the restriction of $\widetilde{\mathcal{P}}$ to the cells associated with $(\mathbf{a},\mathbf{b})$. For each $1\leq i\leq K$, write
@@
Using Equations˜7 and 8, we deduce:
@@
We can decompose the family $\widetilde{\mathcal{P}}\cap R_{i_{0}}=(\mathcal{P}\cap A)\sqcup(\mathcal{P}\cap B)$ into its points above and below $C_{i_{0},i_{0}}$. Since we have $\mathrm{Leb}(A)\leq\Delta x<\delta^{2}/16$, ˜3.2 yields:
@@
and likewise for $B$. ˜3.3 asserts that the sets $\mathcal{P}\cap A$ and $\mathcal{P}\cap B$ consist of i.i.d. uniform points, hence:
@@
The first term of (9) is handled in a similar way. If $i\neq i_{0}$, by ˜3.2:
@@
We can then use ˜3.3 and [S98, Theorem 2] as before:
@@
Finally, putting together Equations˜9, 10 and 11:
@@
for some $c_{\delta}>0$. Therefore, by ˜3.3 and [DZ99, Theorem 1]:
@@
where $\delta>0$ is arbitrary. This concludes the proof of Theorem˜2.1. ∎
Proof of ˜2.2.
We work conditionally given $t^{(n)}$. Then for each $n$, $\tau_{n}$ is a uniform $t^{(n)}$-cyclic permutation. Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$ and $\check{\mathcal{P}}$ be its points outside the diagonal. Define $\check{t}^{(n)}$ as the cycle type $(0,t_{2}^{(n)},t_{3}^{(n)},\dots)$ of size $\check{n}:=n-t_{1}^{(n)}$. Then according to ˜1.3, $\check{\tau}_{n}:={\rm Perm}\left(\check{\mathcal{P}}\right)$ is a uniform $\check{t}^{(n)}$-cyclic permutation. Since fixed points contribute at most $1$ to $\mathrm{LDS}(\tau_{n})$, we can write:
@@
Thus by Theorem˜2.1, for any $\delta>0$ there exists a universal constant $c_{\delta}>0$ such that for any $n$:
@@
and likewise for the lower tail bound. If $\check{n}\to\infty$ in probability then the right-hand side goes to 0 as $n\to\infty$, and this proves convergence in probability. Convergence in $L^{q}$ follows as in the proof of Theorem˜2.1, bounding $\mathrm{LDS}(\tau_{n})$ above by $n-t_{1}^{(n)}+1$ and using the fact that the previous tail inequalities hold a.s. conditionally given $t^{(n)}$. ∎
@@ 4.2 Increasing subsequences
The study of increasing subsequences is subtler than that of decreasing ones in our geometric construction. Indeed, while down-right paths tend to stay away from the diagonal of $0,1]^{2}$, allowing us to use ˜3.3, up-right paths tend to stay very close to it. Consequently, we settle for mere bounds here.
Proof of ˜2.3.
Fix $\delta>0$. Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$ as in ˜1.3.
@@
Using ˜3.2, there exists $c_{\delta}>0$ such that:
@@
Then by [DZ99, Theorem 1] along with ˜3.3, denoting by $L_{m}$ the $\mathrm{LIS}$ of a uniform permutation of size $m$:
@@
since $\frac{1-\delta/2}{K}>\frac{1-\delta}{K-1}$. Finally, using Equation˜12:
@@
For this, simply use ˜3.1 and a rough union bound. Since
@@
Proof of ˜2.4.
The method is the same as in the proof of ˜2.2, using rather the bounds
@@ 5 Proofs of the results on Robinson–Schensted shapes
In [S23], Sjöstrand studied the limit RS shape of so-called locally uniform permutations. The latter are random permutations of the form $\sigma_{n}={\rm Perm}\left(\mathcal{P}_{n}\right)$ where $\mathcal{P}_{n}=\{Z_{1},\dots,Z_{n}\}$ are i.i.d. random points distributed under some absolutely continuous distribution $\rho$. In particular, he proved [S23, Theorem 10.2] that the limit
$\lim_{n\to\infty}\frac{1}{n}\mathrm{LDS}_{r\sqrt{n}}\left(\sigma_{n}\right)$
exists for any $r\geq 0$. It is denoted by $F_{\max}(r)$, where the dependence in $\rho$ is implicitly understood, and is characterized by some variational problem:
$F_{\max}(r)=\sup_{u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)}F_{\rho}(u).$
Precise definitions are postponed to Section˜5.1. The approach in [S23] relies on a shift in point of view: instead of looking for large $r\sqrt{n}$-decreasing subsequences in $\sigma_{n}$, one looks for suitable functions $w_{n}\in\mathcal{U}_{r\sqrt{n}}\left((0,1)^{2}\right)$ that “fit” the point set $\mathcal{P}_{n}$. The quantity $F_{\rho}(u)$ is constructed so that it can be thought of as a “score” associated with $u$, that is, a measurement of how well the function $u$ should fit the decreasing subsequences of $\mathcal{P}_{n}$ after rescaling. [S23, Proposition 4.5] shows that the function $F_{\rho}$ can be used to locally approximate $\mathrm{LDS}_{r\sqrt{n}}\left(\sigma_{n}\right)$; it is also reformulated in handier terms into [S23, Lemma 8.2] which yields a global upper bound, and [S23, Lemma 6.3] which yields a local lower bound.
The difference between our setting and Sjöstrand’s is that the points of our geometric construction $\mathcal{P}$ are not globally independent. Nonetheless, they are locally i.i.d. thanks to ˜3.3, allowing us to use the same techniques as in [S23]. Moreover, the local sampling density is constant in our case ($\rho:=1$), thus greatly simplifying several steps of the proof.
This section is organized as follows. First, in Section˜5.1, we recall the definitions and results from [S23] that will be relevant in our setting. Then, in Section˜5.2, we illustrate them with the case of i.i.d. points sampled uniformly inside an arbitrary domain. Our contribution comes in Section˜5.3, where we use these results to establish Theorem˜2.7, and finally ˜2.8.
5.1 Preliminaries
Figure 4: Left: a doubly increasing function $u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ for some $3\leq r<4$, with its integer level lines. Darker colors indicate greater values of $u$. The set $D(u)$ is highlighted in red. Right: in red, a $3$-decreasing point set $P$. The function $\kappa_{P}$ indeed satisfies $D(\kappa_{P})=P$. Its values range from 0 to $3$, with lighter colors indicating lower values and darker colors indicating higher values.
Let $\Omega$ be an open subset of $\mathbb{R}^{2}$. If $u:\Omega\to\mathbb{R}$ is Lebesgue-measurable, write $\lVert u\rVert_{1,\Omega}:=\int_{\Omega}\left|u\right|d\mathrm{Leb}$ for its $L^{1}$ norm. Recall that we use the partial order on $\mathbb{R}^{2}$ defined by:
$(x,y)$ $\leq(x^{\prime},y^{\prime})\quad\text{ if and only if }\quad x\leq x^{\prime}\text{ and }y\leq y^{\prime}$.
Then, we call $u$ doubly increasing if $u(x,y)\leq u(x^{\prime},y^{\prime})$ whenever $(x,y)\leq(x^{\prime},y^{\prime})$. Denote by $\mathcal{U}(\Omega)$ the set of doubly increasing functions on $\Omega$, and by $\mathcal{U}_{h,r}(\Omega)$ its subset consisting of functions with values in $[h,h+r]$. For any $u\in\mathcal{U}(\Omega)$, define
@@ 5.1 Preliminaries
In words, $D(u)$ is the set of “south-west corners” in the integer level lines of $u$. Also, if $P$ is a finite subset of $\Omega$, define $\kappa_{P}:(x,y)\in\Omega\mapsto\mathrm{LIS}\left(\big.P\cap\left((-\infty,x]\times(-\infty,y]\right)\right)$. See Figure˜4 for an illustration.
It is easy to check that if $u\in\mathcal{U}_{0,r}(\Omega)$ then $D(u)$ is an $\lfloor r\rfloor$-decreasing subset of $\Omega$, see [S23, Lemma 6.2]. Reciprocally, one can also check that if $P$ is a $k$-decreasing subset of $\Omega$ then $\kappa_{P}$ is in $\mathcal{U}_{0,k}(\Omega)$ and satisfies $D(\kappa_{P})=P$, see again [S23, Lemma 6.2]. Now, if $\mathcal{P}$ is a finite point set in $\Omega$, recall that $\mathrm{LDS}_{k}(\mathcal{P})$ can be expressed as the maximum of $\left|\mathcal{P}\cap P\right|$ where $P$ ranges over $k$-decreasing subsets of $\Omega$. Subsequently, by [S23, Lemma 6.2]:
$\mathrm{LDS}_{k}(\mathcal{P})=\max_{u\in\mathcal{U}_{0,k}(\Omega)}\left|D(u)\cap\mathcal{P}\right|.$
We shall use this equation after rescaling, that is: